### v1

# Molecular Communication Channel as a Physical Reservoir Computer

Abstract
Molecular Communication (MC) channels inherently possess significant memory and nonlinear dynamics due to diffusion and receptor kinetics, often posing challenges for reliable data transmission. This work reconceptualizes these intrinsic properties as computational resources by framing a canonical point-to-point MC channel, consisting of ligand diffusion and reversible ligand-receptor binding at a spherical receiver, as a physical reservoir computer (PRC). We utilize the time-varying fraction of bound receptors as the reservoir’s internal state, employing time-multiplexing to generate high-dimensional virtual nodes without explicit recurrence. Only a linear readout layer is trained via ridge regression. Through deterministic mean-field modeling and particle-based spatial stochastic simulations, we demonstrate the MC system’s capability for complex temporal processing by successfully performing next-step prediction on standard chaotic time-series benchmarks (Mackey-Glass and NARMA10). Performance, quantified by Normalized Root Mean Square Error (NRMSE), exhibits a non-monotonic dependence on key system parameters (receptor kinetic rates, diffusion coefficient, transmitter-receiver distance), revealing optimal operational regimes. These findings validate the potential of using MC channel as effective and low-complexity computational substrate.
Index Terms:
Reservoir Computing, Molecular Communications, Ligand-Receptor Interactions
I Introduction
Reservoir Computing (RC), introduced independently as the echo‑state network (ESN) by Jaeger [1] and the liquid‑state machine (LSM) by Maass et al. [2], is a class of recurrent neural network models in which an untrained, nonlinear dynamical system (the reservoir) projects an input into a high‑dimensional state space, while only a linear readout is trained, typically via ridge regression. This separation of representation and learning obviates back‑propagation through time, providing fast, data‑efficient training while retaining the rich temporal processing power of recurrent dynamics. As a result, RC has proven highly effective for time-series prediction tasks [3].
Physical Reservoir Computing (PRC) generalizes this principle by exploiting the native dynamics of physical substrates, e.g., photonic [4], electronic [5], spintronic [6], and mechanical systems [7], as reservoirs, often in delay‑based architectures that implements large virtual networks through time-multiplexing of a single nonlinear node [8, 9]. Such implementations greatly reduce hardware complexity while achieving performance on par with digital RC, and have catalyzed interest in neuromorphic edge computing [10] and embodied intelligence [11].
Recent work has begun to evaluate biological systems as reservoirs, capitalizing on their inherently rich and adaptive dynamics. Pioneering experiments with neuron cultures [12], brain organoids [13], bacteria [14], and plants [15] demonstrate that living systems can be co-opted or integrated into artificial computing frameworks to execute externally defined computational tasks. Likewise, chemical reaction networks, which underlie many fundamental biochemical processes, have been investigated for RC, demonstrating the potential to exploit complex molecular interactions for sophisticated information processing beyond their native biological context [16].
Central to the exceptional information processing capabilities of biological systems is biochemical signaling, which enables coordination and adaptation across multiple scales, critical for their survival [17]. Building on these naturally evolved mechanisms, researchers have long studied engineered molecular communication (MC) systems as alternative to conventional electromagnetic (EM) communications to enable networks of micro/nanoscale devices via the exchange of precisely controlled chemical signals [18]. This line of work has fostered the vision of the Internet of Bio-Nano Things (IoBNT), in which biological entities and artificial micro/nanodevices interface seamlessly for next-generation biomedical applications [19]. Extensive theoretical research and recent experimental work in this field have particularly focused on overcoming the unique challenges of MC channels that are different than EM communications, including significant channel memory effects (inter-symbol interference) and nonlinear receiver responses (involving receptor saturation) in addition to biochemical noise [20, 21]. These complexities turned out to be particularly demanding for small, resource-limited devices, yet they also hint at untapped computational potential that biological systems may be inherently exploiting.
In this paper, we reconceptualize MC channel as a physical reservoir computer, revealing its potential for information processing and implementing machine learning tasks. Specifically, we study a simple MC channel model, which is adequately representative of both biological and synthetic MC systems, consisting of a point transmitter and a spherical receiver whose surface is functionalized with ligand receptors. By adopting a time-multiplexed RC framework, we leverage the inherent channel memory arising from diffusion and the nonlinear ligand-receptor interactions at the receiver, both of which are critical for computational performance. Through numerical analysis and particle-based spatial stochastic simulations, we demonstrate that MC channels can successfully perform chaotic time-series prediction tasks with minimal training, as benchmarked by standard RC evaluations on NARMA10 and Mackey-Glass time-series. We systematically investigate how various MC system parameters, including the number of transmitted molecules, transmitter-receiver distance, diffusion coefficient, and ligand-receptor kinetic rates, affect computing performance.
Our results provide evidence that even simple MC channels can be harnessed for non-trivial time-series tasks by virtue of their intrinsic diffusion-driven memory and receptor-mediated nonlinearities. Beyond contributing to the fundamental understanding of information-processing capabilities of biological systems at the elemental level of their communication links, this computing perspective on MC channels can open new avenues for bio-inspired and bio-integrated computing architectures. Potential applications range from early disease diagnosis analyzing time-varying biomarker profiles to channel equalization in artificial MC networks and IoBNT [22]. By demonstrating how the channel’s inherent dynamics can be systematically exploited for information processing, this study lays the groundwork for MC systems to serve not only as physical channels for communication but also as powerful computational substrates in their own right.
The paper is organized as follows. Section II reviews molecular communication and physical reservoir computing. Section III formalizes the molecular‑communication channel as a physical reservoir computer and describes the modeling and training methodology. Section IV presents deterministic and stochastic simulation results with parameter studies. Section V concludes and outlines future directions.
II Background
II-A Molecular Communications
Bio-inspired MC emerged as a promising approach to realize networks of micro/nanoscale devices, such as biosensors and microrobots, and to seamlessly interface these engineered networks with biological systems. Such integration underpins the vision of the IoBNT, where MC mediates collaborative tasks such as targeted drug delivery and continuous health monitoring [23]. By exploiting mechanisms similar to cellular signaling, MC provides bio-compatible, low-energy connectivity in environments where conventional EM methods become impractical [24].
From early theoretical studies on capacity and error rates, the MC field has advanced to increasingly sophisticated experiments [21, 25, 26, 27]. Researchers have identified unique challenges that distinguish MC from EM communications, including severe inter-symbol interference (ISI) due to slow diffusion, and nonlinearity plus memory effects from stochastic receptor binding and other reaction kinetics. Practical MC transceivers at the microscale also face tight constraints in energy, size, and computational resources, motivating new strategies in encoding, detection, and synchronization [20]. To address these issues, data-driven and machine learning methods have emerged, with neural networks learning the input-output mapping of MC channels directly from data and thus coping better with time-varying conditions and nonlinear dynamics [28].
In parallel, an expanding line of research investigates MC for in situ artificial intelligence (AI) with living cells and engineered biomolecular processes, commonly referred to as wetware AI. Two complementary strategies have emerged: top-down approaches that leverage existing cellular functionalities, and bottom-up approaches that build computing frameworks from fundamental chemical reactions. In the former category, wet-neuromorphic computing was introduced by repurposing gene regulatory networks (GRNs) for pattern recognition [29], while Wet TinyML concept applied GRN-based methods for energy-efficient neural-like computations within living cells [30]. Another study integrated biologically plausible spiking neurons into reservoir computing, using MC-based connectivity for distributed processing among cells [31]. A unifying framework for Molecular Machine Learning (MML) positioned microbial consortia in addition to GNRs within individual cells as native substrates for AI, with MC serving as the core mechanism for data exchange and biochemical learning [32].
By contrast, bottom-up efforts explicitly engineer MC as the foundation for novel neural network architectures. Nano-scale computing units and Molecular Nano Neural Networks (M3N) have been proposed, wherein diffusion-based molecular signaling implements neural-layer functions for in-body intelligence [33]. Another line of work introduced a broader vision of chemical AI, illustrating how reaction networks, synthetic biology, and nanofluidic systems, underpinned by molecular signaling, could equip synthetic cells with self-contained, biocompatible computation [34].
Altogether, these works highlight the synergy between biological information processing, MC, and ML in paving the way for low-power and adaptive intelligence at micro- and nanoscales and within biological contexts. Here, we introduce a new perspective by reframing an elemental point-to-point channel, comprising diffusion and receptor-ligand interactions, as a physical reservoir computer. Unlike previous studies relying on genetic or multi-cellular networks and specialized biochemical circuits, our focus is on systematically exploiting the channel’s intrinsic memory and nonlinearity to perform non-trivial learning tasks with minimal training overhead. Our findings reveal new pathways for leveraging MC channels as both a communication medium and a computational resource, thereby extending earlier demonstrations of wetware AI.
Figure 1: (a) RC framework for a multiclass classification task, (b) An example of physical RC for binary classification: the wave-interference patterns produce a nonlinear mapping of the data, which is then trained for classification [35], (c) Delay-based RC framework with a singular nonlinear node for multiclass classification task, (d) An end-to-end MC system as a reservoir computer with virtual nodes for time-series prediction task.
II-B Physical Reservoir Computing
In a standard RC scheme, the input vector $\mathbf{u}(t)$ drives the reservoir state $\mathbf{x}(t)\in\mathbb{R}^{N}$ through
$\mathbf{x}(t+1)=f\!\bigl{(}\mathbf{A}\,\mathbf{x}(t)+\mathbf{W}_{\mathrm{in}}\mathbf{u}(t)\bigr{)},$ (1)
where $\mathbf{A}\in\mathbb{R}^{N\times N}$ is a fixed (typically random) recurrent weight matrix, $\mathbf{W}_{\mathrm{in}}$ projects the input into the reservoir, and $f(\cdot)$ is an element‑wise nonlinearity (Fig. 1(a)) [36]. The readout is obtained by a linear map
$\mathbf{y}(t)=\mathbf{W}_{\mathrm{out}}\mathbf{x}(t),$ (2)
with $\mathbf{W}_{\mathrm{out}}$ trained, often by ridge regression, to minimize the discrepancy between $\mathbf{y}(t)$ and a target signal. Because the internal weights $\mathbf{A}$ remain fixed, RC sidesteps the heavy cost of back‑propagating gradients through deep recurrent structures, a feature that makes it well‑suited to hardware realizations.
PRC replaces the software‑defined reservoir with the intrinsic dynamics of a physical substrate. Phenomena such as optical feedback, spin‑wave interactions, analogue electronic circuits, and even the interference patterns of water waves in a bucket implement the operations $\mathbf{A}\mathbf{x}(t)$ and $f(\cdot)$ in real time, while external sensors capture the instantaneous state $\mathbf{x}(t)$ (Fig. 1(b)) [9, 35]. Leveraging the medium’s native parallelism, bandwidth, and energy efficiency, PRC offers a neuromorphic alternative to digital simulations.
Time‑multiplexed reservoirs. High‑dimensional states can be generated with minimal hardware by time‑multiplexing a single nonlinear element (Fig. 1(c)) [8]. During each clock period of length $T$, the input segment is multiplied by a masking sequence $\mathbf{m}$, and the element’s output is sampled at $M$ equidistant instants. The resulting vector
$\mathbf{x}(t)=\bigl{[}x_{1}(t),x_{2}(t),\dots,x_{M}(t)\bigr{]}^{\!\top}$
acts as an $M$‑node virtual reservoir whose evolution from one period to the next is governed by the masked input and the device’s nonlinear response [37].
A widely adopted form of time‑multiplexing employs an external feedback loop of delay $\tau$ around a single nonlinear node (e.g., a semiconductor laser or an analogue circuit). Partitioning the delay into $M$ sub‑intervals of length $\Delta=\tau/M$ yields $M$ virtual nodes [9, 38]. Sampling the output every $\Delta$ seconds constructs $\mathbf{x}(t)$, while the recirculating signal supplies fading memory that underpins tasks such as speech recognition and chaotic time‑series forecasting. Additional delay lines or delayed taps at the input or readout can further enrich the dynamics or extend the memory horizon without increasing the number of physical components [39].
III Molecular Communication Channel as a Physical Reservoir
We consider a single-link MC system consisting of a point transmitter and a spherical receiver, as illustrated in Fig. 1(d). The transmitter release a prescribed number of ligands (molecules) instantaneously at prearranged times into a fluidic medium, where they propagate by free diffusion. A fraction of these ligands eventually arrives near the receiver and can bind reversibly to receptors on its surface under given reaction kinetics. Below, we describe how this widely studied MC system model can be reframed as a physical reservoir computer by regarding the time-varying fraction of bound receptors as the high-dimensional dynamical state.
Let the transmitter release $I(n)$ molecules at each symbol time $nT$, where $n\in\{0,1,2,\dots\}$, and $T$ is the symbol interval. For simplicity, we assume the receiver is at distance $x_{R}\gg r_{\mathrm{rx}}$ from the transmitter at the origin, such that the receiver’s finite radius $r_{\mathrm{rx}}$ can be neglected in the channel impulse response. Under three-dimensional (3D) free diffusion with diffusion coefficient $D$, the ligand concentration $c(\mathbf{x},t)$ obeys Fick’s second law:
$\frac{\partial c(\mathbf{x},t)}{\partial t}\;=\;D\,\nabla^{2}c(\mathbf{x},t),$ (3)
with the far-field boundary condition $c(\mathbf{x},t)\to 0$ as $\|\mathbf{x}\|\to\infty$. The impulse response at distance $x_{0}$ for a single pulse released at $t=0$ is the 3D diffusion kernel
$h(t)\;=\;\frac{1}{(4\pi Dt)^{3/2}}\exp\!\Bigl{(}-\frac{x_{0}^{2}}{4Dt}\Bigr{)}\quad(t>0).$ (4)
Hence, when the transmitter emits $I(n)$ molecules at times $nT$, the local ligand concentration at the receiver (treated as a point in the far-field approximation) is given by
$c_{T}(t)\;=\;\sum_{n=0}^{\infty}I(n)\,h\bigl{(}t-nT\bigr{)}\,u\bigl{(}t-nT\bigr{)},$ (5)
where $u(\cdot)$ is the unit step function enforcing causality. Since $h(t)$ has a infinitely long tail (though it decays rapidly), successive transmissions introduce ISI. For practical numerical simulations, this tail is often truncated to a finite channel memory length [40].
Figure 2: (a) An instance from particle-based stochastic simulations of the MC channel with Smoldyn. Receptor occupation ratios for numerical and stochastic simulations of the MC system for (b) Mackey-Glass series input, (c) NARMA10 Series input, (d) Mackey-Glass series input, post filtering (smoothing), (e) NARMA10 series input, post filtering.
The spherical receiver is equipped with $N_{R}$ identical, monovalent ligand receptors. Each receptor can be either unbound or bound to one ligand, with reversible reaction:
$L+R\overset{k_{\text{on}}}{\underset{k_{\text{off}}}{\longleftrightarrow}}LR,$
where $L$ is the free ligand, $R$ is an unbound receptor, and $LR$ is a ligand-receptor complex, i.e., bound receptor. At the mean-field level, and assuming the local free ligand concentration is not significantly depleted by bound receptors, the time evolution of the mean fraction of bound receptors, $b(t)=n_{b}(t)/N_{R}$, can be approximated by
$\frac{d\,b(t)}{dt}\;=\;k_{\mathrm{on}}\,c_{T}(t)\,\bigl{[}1-b(t)\bigr{]}\;-\;k_{\mathrm{off}}\,b(t),$ (6)
where $k_{\mathrm{on}}$ and $k_{\mathrm{off}}$ are the forward (binding) and reverse (unbinding) rate constants, respectively, and $c_{T}(t)$ is the local ligand concentration at the receiver.
Although the exact time-varying variance of $b(t)$ under such nonstationary input is typically not available in closed form (due to correlated ligand arrivals and receptor occupancy fluctuations), we use particle-based spatial stochastic simulations (with Smoldyn (Fig. 2) [41]) in Section IV to account for this stochasticity. For the present analytical treatment, (6) provides a deterministic approximation to the mean behavior of the ligand-receptor interactions.
[…]
In RC, one aims to transform an input time series $\{u(n)\}$ into a high-dimensional dynamical state $\mathbf{x}_{n}$ through a fixed nonlinear system (the reservoir), and then train only a linear readout for the target computational task. Here, we map the input $u(n)$ to the number of released molecules $I(n)$ at each symbol time $nT$:
$I(n)\;=\;\alpha\,u(n),$ (7)
where $\alpha$ is a scaling factor. The channel then imposes a complex, memory-rich temporal response on these inputs (due to both diffusion and receptor binding).
To extract a sufficiently high-dimensional state from the scalar output $b(t)$, we perform time-multiplexing. That is, within each interval $[nT,\,(n+1)T]$, we sample $b(t)$ at $M$ distinct time offsets $\{\tau_{1},\tau_{2},\dots,\tau_{M}\}$:
$\mathbf{x}_{n}\;=\;\bigl{[}b\bigl{(}nT+\tau_{1}\bigr{)},\;b\bigl{(}nT+\tau_{2}\bigr{)},\;\dots,\;b\bigl{(}nT+\tau_{M}\bigr{)}\bigr{]}^{\top}.$ (8)
This sampling effectively creates $M$ virtual nodes per symbol. Unlike a typical optoelectronic delay-based reservoir, however, we do not require a separate external feedback loop; rather, the channel’s intrinsic ISI provides the dynamical memory. The “masking” function typically used in conventional time-multiplexed RC is replaced here by the MC channel impulse response, which shapes the input in time to eventually generate diverse node responses upon sampling.
We train a linear readout to map from the reservoir state $\mathbf{x}_{n}$ to the desired output $y_{n}$ (e.g., a time-series prediction). Let $\mathbf{W}_{\mathrm{out}}$ be the $M$-dimensional readout weight vector, then
$\hat{y}_{n}\;=\;\mathbf{W}_{\mathrm{out}}^{\!\top}\,\mathbf{x}_{n}.$ (9)
With $N_{\mathrm{train}}$ training samples $\{(\mathbf{x}_{n},\,y_{n})\}$, we solve:
$\min_{\mathbf{W}_{\mathrm{out}}}\Bigl{\|}\mathbf{y}\;-\;\mathbf{X}\,\mathbf{W}_{\mathrm{out}}\Bigr{\|}^{2}\;+\;\lambda\,\|\mathbf{W}_{\mathrm{out}}\|^{2},$ (10)
where $\mathbf{X}$ is the $N_{\mathrm{train}}\!\times\!M$ matrix whose $n$-th row is $\mathbf{x}_{n}^{\top}$, $\mathbf{y}$ is the vector of target outputs $y_{n}$, and $\lambda$ is a regularization parameter (ridge regression). The closed-form solution is then given by the Moore–Penrose pseudoinverse:
$\mathbf{W}_{\mathrm{out}}\;=\;\bigl{(}\mathbf{X}^{\top}\,\mathbf{X}\;+\;\lambda\,\mathbf{I}\bigr{)}^{\!-1}\,\mathbf{X}^{\top}\,\mathbf{y}.$ (11)
Hence, only $\mathbf{W}_{\mathrm{out}}$ is optimized, and the underlying channel and binding dynamics remain unmodified.
IV Results
In this section, we present the results of the performance evaluation for the MC reservoir computer using two complementary approaches. First, deterministic analyses were carried out in MATLAB using the mean‑field kinetic model (Eq. (6)) with the default simulation settings listed in Table I. Second, to capture the stochasticity of the MC channel, we performed particle‑based spatial stochastic simulations in Smoldyn [41]. By comparing the outcomes of the deterministic and stochastic simulations, we both confirm the validity of the approximate deterministic model and quantify the effects of intrinsic noise on RC performance.
TABLE I: Default Values of Simulation and RC Parameters
Parameter Value
Receptor binding constant $k_{\text{on}}$ $6.022\times 10^{8}\,\mathrm{M^{-1}s^{-1}}$
Receptor unbinding constant $k_{\text{off}}$ $1\,\mathrm{s^{-1}}$
Maximum # released molecules $N_{\max}$ $3000$
Minimum # released molecules $N_{\min}$ $100$
Symbol duration $T$ $1\,\mathrm{s}$
Mackey–Glass prediction length $\Delta$ $6$ steps
Diffusion coefficient $D$ $1\times 10^{-11}\,\mathrm{m^{2}s^{-1}}$
Tx–Rx distance $x_{0}$ $10\,\upmu\mathrm{m}$
Reservoir nodes $M$ $100$
Numerical (deterministic) soln. timestep $\Delta t_{d}$ $1\,\mathrm{ms}$
Stochastic simulation timestep $\Delta t_{s}$ $10\,\mathrm{ms}$
Stochastic simulation receiver radius $R$ $3\,\upmu\mathrm{m}$
Training sequence length $2000$
Test sequence length $2000$
IV-A Chaotic Time-Series Prediction
We benchmark our MC reservoir computer on two widely used chaotic time‑series prediction tasks: Mackey–Glass and NARMA10 [42]. These tasks have become standards for assessing both algorithmic and physical RC architectures because they combine complex temporal dependencies, strong nonlinearities, and a requirement for long‑range memory. The goal is to learn a mapping from the present (and past) states of the time series to future values, using only a linear readout layer trained on the reservoir states.
Performance Metric: We quantify the predictive accuracy of the reservoir using the normalized root-mean-square error (NRMSE),
$\text{NRMSE}\;=\;\sqrt{\frac{\sum_{k=1}^{N_{\mathrm{test}}}[\,y_{k}-\hat{y}_{k}\,]^{2}}{\sum_{k=1}^{N_{\mathrm{test}}}[\,y_{k}-\bar{y}\,]^{2}}},$
where $y_{k}$ is the true target value at time $k$, $\hat{y}_{k}$ is the reservoir’s predicted value, $\bar{y}$ is the mean of all true target values, and $N_{\mathrm{test}}$ is the number of test samples. An NRMSE closer to 00 indicates a more accurate prediction.
Figure 4: Effects of MC reservoir parameters on computational performance in terms of NMRSE, (a-c) for Mackey-Glass series, and (d-f) for NARMA10 series prediction tasks.
IV-A1 Mackey-Glass Series
The Mackey–Glass system is described by a time-delayed differential equation originally introduced to model physiological control processes such as blood production:
$\frac{dx(t)}{dt}=\beta\,\frac{x(t-\tau)}{1+x(t-\tau)^{n}}-\gamma\,x(t),$ (12)
where $\beta$, $\gamma$, $n$, and $\tau$ are positive constants governing the nonlinear feedback loop. For the canonical parameter set $\beta=0.2$, $\gamma=0.1$, $n=10$, and $\tau=17$, the system exhibits chaotic behavior with a fractal attractor.
We generate a univariate time series $\{x_{k}\}$ by numerically integrating the Mackey–Glass equation using a suitable step size (e.g., $0.1$), then sampling at discrete time points $k$. The reservoir input at time step $k$ is set to $u(k)=x_{k}$, and the task is to predict $x_{k+\Delta}$, i.e., the value of the series $\Delta$ steps into the future.
IV-A2 NARMA10 Series
A complementary benchmark is the NARMA10 (tenth‑order Non‑linear Auto‑Regressive Moving‑Average) series, defined by
$y(k+1)$ $=0.3\,y(k)\;+\;$ (13)
$0.05\,y(k)\sum_{i=0}^{9}y(k-i)+1.5\,u(k-9)u(k)+0.1,$
where $y(k)$ is the output sequence and $u(k)$ is a scalar driving input (often i.i.d. uniform in [0, 0.5]00.5[0,\,0.5][ 0 , 0.5 ]). In this one-step-ahead prediction task, the reservoir receives $u(k)$ at time $k$ and must produce $\hat{y}(k+1)$ that approximates the true next value $y(k+1)$. The main challenge is that $y(k+1)$ depends on its previous ten outputs in a highly nonlinear manner, requiring the reservoir to capture temporal dependencies spanning at least ten steps.
IV-A3 Discussion
Fig. 3 summarizes the MC reservoir’s predictive performance on these tasks under different simulation conditions (rows) and time horizons (columns). The leftmost column illustrates the reservoir’s one-step-ahead forecasts for the NARMA10 series; the remaining three columns correspond to Mackey–Glass predictions at horizons of 1, 10, and 20 steps, respectively. Within each panel, the true time series (blue) is overlaid with the predicted values (red).
Rows (a–d) in Fig. 3 show the deterministic mean-field simulation results. Rows (e–h) show particle-based stochastic simulations that capture molecular noise. We observe that although the more realistic stochastic simulations introduce additional fluctuations in the reservoir node states, the overall predictive accuracy remains quite comparable to the deterministic case, indicating that the approximate mean-field analysis is valid and that the fluctuations do not severely degrade performance.
Focusing on the Mackey–Glass predictions (columns 2–4), we see that the reservoir accurately tracks the chaotic trajectory for modest horizons (e.g., $\Delta=1$ or $\Delta=10$), but it can still capture essential dynamics even at $\Delta=20$. Unsurprisingly, the error grows at larger horizons, yet the predictions remain qualitatively coherent with the true signal. This capability underscores the MC reservoir’s capacity for handling long-range memory in a strongly nonlinear system.
In the leftmost column, the reservoir’s one-step predictions for NARMA10 closely match the target sequence. Although NARMA10 typically requires the memory of up to ten past outputs, the reservoir’s nonlinear fading dynamics prove sufficient to encode that history and generate accurate forecasts.
IV-B Optimization of the Reservoir
IV-B1 Effect on MC System Parameters on Performance
The computational efficacy of the proposed molecular reservoir computer is intricately related to the biophysical parameters that govern the MC channel. Specifically, the dynamics of ligand-receptor binding, molecular propagation, and symbol encoding collectively define the temporal complexity and nonlinearity accessible to the reservoir. To systematically evaluate these dependencies, we conducted a parametric study involving some key variables: the ligand-receptor binding and unbinding rate parameters governing the receiver response {$k_{on}$, $k_{off}$}, the symbol interval and input amplitude governing transmitted signal{$N_{max}$, $T$}, and the propagation channel parameters such as diffusion coefficient (D) and transmitter-receiver distance $x_{0}$. The effects of changing these parameters in a physically plausible range are presented in Fig. 4 through the metric of NRMSE.
Our findings reveal a nonmonotonic relationship between parameter settings and predictive performance, as measured by the NRMSE. Specifically, the interplay between binding dynamics (i.e., the ratio $k_{off}$/$k_{on}$), inter-symbol duration, diffusion rate, and transmitter-receiver distance plays a critical role in determining the channel’s computational effectiveness. Low $k_{off}$/$k_{on}$ ratios, short inter-symbol durations, and high diffusion coefficients—especially when combined with shorter distances—lead to receiver saturation. In these scenarios, the occupation ratio of the receptor remains consistently high, nearing full saturation with minimal temporal variation. As a result, the system exhibits reduced dynamical richness and a diminished capacity to encode time-varying input features, which leads to an increase in NRMSE and degraded predictive accuracy. This effect is particularly evident in high-density ligand environments where excessive receptor occupation eliminates the temporal diversity necessary for reservoir computation.
Conversely, conditions characterized by high $k_{off}$/$k_{on}$ ratios or insufficient diffusion coefficients impede the ability of ligands to effectively bind and form receptor-ligand complexes. In such settings, the MC channel fails to facilitate reliable signal transmission, leading to sparse or inconsistent occupation patterns. This not only reduces the input-driven modulation of the reservoir state but also undermines the system’s ability to perform the nonlinear transformation essential for projecting input signals into a higher-dimensional computational space. Moreover, the consistent trends in prediction performance observed across both the NARMA10 and Mackey–Glass benchmarks reinforce the generalizability of these parameter–performance relationships, demonstrating that the interplay between binding kinetics and diffusion dynamics fundamentally governs the temporal richness and nonlinearity of the reservoir, and thus provides clear guidelines for optimizing MC‑based reservoir computers under diverse biochemical conditions.
Figure 5: Effect of the moving average filter window length on the NRMSE for stochastic simulations.
IV-B2 Post-Process Filtering
Due to the inherent stochasticity of molecular diffusion and receptor-ligand interactions, especially in low-concentration regimes, the raw output of the molecular reservoir exhibits noise-driven fluctuations, as observed in Fig. 2. These fluctuations, though biologically realistic, can obfuscate the underlying information-carrying trends. To address this, we applied a moving average filter to the received signal of receptor occupation $b(t)$, effectively attenuating high-frequency noise components while preserving meaningful temporal dynamics. The entire received signal sequence is filtered, the effects of which on the prediction task are shown in Fig. 3. The post-processed received signals are observed to display outputs similar to those of the estimation layer to deterministic simulations, enabling more robust performance during inference. The change in error with respect to the length of the filtering window is presented in Fig. 5, the filter improves performance until a certain point, after which the loss of information becomes too great and leads to an increase in error.
V Conclusion
In this work, we have demonstrated, through cross‑validated deterministic and particle‑based stochastic simulations, that a canonical point‑to‑point MC channel naturally realizes a physical reservoir computer. By interpreting the time‑multiplexed receptor occupancy as virtual nodes, the system reliably forecasts Mackey–Glass and NARMA10 time series with low normalized error, despite substantial molecular noise. A comprehensive parameter sweep analysis revealed that computational power of the MC reservoir hinges on a delicate balance between several system parameters, including receptor binding kinetics, diffusion coefficients, and symbol intervals, all of which influence the memory and nonlinearity essential for effective reservoir computing. This analysis provides the initial guidelines for rationally optimizing MC systems for wetware AI. Future work will focus on extending the system to MC networks and multiple ligand-receptor pairs, and on implementing the proposed physical reservoir computer in experimental MC platforms.